OpenStudy (anonymous):

Can somebody explain how to simplify this expression? Negative 2y to the negative 1 power, all inside parentheses, with a negative 2 exponent outside the parentheses.

OpenStudy (amistre64):

using math notation for staters

OpenStudy (amistre64):

\[(-2y^{-1})^{-2}\]

OpenStudy (amistre64):

this?

OpenStudy (amistre64):

\[-2^{-2} y^2\] \[-\frac{y^2}{4}\] maybe?

OpenStudy (radar):

Is it still a minus (if it was inside the paren)?

OpenStudy (anonymous):

Yes, you set it up right, amistre.

OpenStudy (radar):

I suspect the answer may be \[y ^{2}\over 4\]

OpenStudy (radar):

Maybe amisstrte64 will come back and verify his answer.

OpenStudy (anonymous):

I think it is 4y^2

OpenStudy (anonymous):

...because it looks like this is the expression... (-2y^-1)^-2 Multiplying exponents (-2y)^2 Expanding brackets 4y^2

OpenStudy (radar):

I went this direction, but I do see your logic. I may have violated order of operations:\[(-2y ^{-1})^{-2}\] to:\[1\over (-2y ^{-1})^{2}\]then to:\[1\over 4 y ^{-2}\] finally\[y ^{2}\over 4\]

OpenStudy (anonymous):

\[(-2y^{-1})^{-2}=(-2)^{-2}y^2=\frac{y^2}{(-2)^2}=\frac{y^2}{4}\]\]

OpenStudy (radar):

Help me out here satellite73

OpenStudy (anonymous):

if that was the problem to beginwith

OpenStudy (anonymous):

we get \[\frac{y^2}{4}\] yes?

OpenStudy (anonymous):

hello radar! here to celebrate amistre's 1000 medal

OpenStudy (radar):

that is what i got, amistre got a -y^2/4

OpenStudy (anonymous):

hope it is soon cause i got to run

OpenStudy (radar):

before you run look at gianfranco solution above, what is wrong with that approach beside getting a different answer???

OpenStudy (anonymous):

well if the question is as he wrote it, you have to raise (-2) to the power of -2, not the same as \[-2^{-2}\]

OpenStudy (anonymous):

you should get a 4 in the denominator

OpenStudy (anonymous):

gianfraco was acting as if it was \[((-2y)^{-1})^{-2}\]

OpenStudy (anonymous):

but that is now how i read the problem. i read it as only the y being raised to the power of -1. if there were no parentheses that is what it means

OpenStudy (anonymous):

i read it the way you did

OpenStudy (radar):

I can see that that would make a difference.

OpenStudy (anonymous):

of course. like the difference between \[(2\times 3)^2\] and \[2\times 3^2\]

OpenStudy (radar):

Understand. Thanks for clearing up a few things.

OpenStudy (anonymous):

O never mind everybody..I think I've figured it out.

OpenStudy (radar):

I was hoping you would